OpenStudy (anonymous):

Find the surface area of a cylinder with d = 12 in: h = 14 in. A. 113.04 sq. in. B. 640.56 sq. in. C. 527.52 sq. in. D. 753.6 sq. in.

OpenStudy (jdoe0001):

\(\bf \textit{total surface area of a cylinder}=2\pi r(h+r) \\ \quad \\ r=radius=\frac{diameter}{2}\qquad \qquad h=height\)

OpenStudy (anonymous):

37.68 + 20 = 58.68 (:

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but its not my choices

OpenStudy (anonymous):

yeah I think that person gave the wrong formula, do you know it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I got it! the answer is D

OpenStudy (anonymous):

thanks nina

OpenStudy (anonymous):

The formula is SA = (2 x PI x r x H) + (2 x PI x (r x r). So when you calculate it you get 753.6 google is all you need if you don't have a formula haha and you're welcome ^.^